Can someone help me solve this?
Solve the following.
sqrt(x+4)=x+4

Answer by edjones(8007)

sqrt(x+4)=x+4
x+4=(x+4)^2 square both sides.
x+4=x^2+8x+16
x^2+7x+12=0 subtract L side from R.
(x+3)(x+4)=0
x=-3, x=-4
.
Ed